### v1

Abstract
We study the trade-off between (average) spread and width in tree decompositions, answering several questions from Wood [arXiv:2509.01140]. The spread of a vertex $v$ in a tree decomposition is the number of bags that contain $v$. Wood asked for which $c>0$, there exists $c^{\prime}$ such that each graph $G$ has a tree decomposition of width $c\operatorname{tw}(G)$ in which each vertex $v$ has spread at most $c^{\prime}(d(v)+1)$. We show that $c\geq 2$ is necessary and that $c>3$ is sufficient. Moreover, we answer a second question fully by showing that near-optimal average spread can be achieved simultaneously with width $O(\operatorname{tw}(G))$.
keywords:
Tree decomposition, spread, domino treewidth
category:
[…]
1 Introduction
Tree decompositions have been widely studied in both structural and algorithmic contexts, for example due to their importance to graph minor theory [21] and the fact that many problems can be solved in linear time on graphs of bounded treewidth [2, 10].
The main criterion of ‘optimality’ for a tree decomposition is usually its width, or, equivalently, the size of the largest bag in the tree decomposition. Recently, Wood [24] proved several results concerning tree decompositions which have additional nice properties, while not sacrificing too much of the width. Our work continues on this, answering several of the problems posed by Wood.
Besides the width, our first parameter of interest is the spread. The spread of a vertex $v$ in a tree-decomposition $(T,(B_{x})_{x\in V(T)})$ is the number of bags $v$ is in: $|\{x\in V(T):v\in B_{x}\}|$. Since all neighbours of a vertex $v$ must appear together with $v$ in some bag, when the width is held constant, the spread of $v$ must necessarily grow with its degree.
Wood [24] posed the following problem.
[…]
What is the infimum of $c$ such that for some $c^{\prime}$, every graph $G$ with treewidth $k$ has a tree-decomposition with width at most $(c+o(1))k$, in which each vertex $v\in V(G)$ has spread at most $c^{\prime}(d(v)+1)$?
Wood [24] already proved that the answer is at most 14 (with $c^{\prime}=1$).
Theorem 1.1 (Theorem 2 in [24]).
Every graph $G$ with treewidth $k$ has a tree-decomposition with width at most $14k+13$, such that each vertex $v\in V(G)$ has spread at most $d(v)+1$.
We improve the lower and upper bound for Problem 1 with the following two results, which show that the value of $c$ of Problem 1 is in the interval $[2,3]$.
Theorem 1.2.
For each $c>3$, there is a constant $c^{\prime}>0$ such that every graph $G$ with treewidth $k$ has a tree-decomposition with width at most $ck$, in which each vertex $v\in V(G)$ has spread at most $c^{\prime}(d(v)+1)$
Theorem 1.3.
For each $c<2$ and for each $c^{\prime}>0$, there is a graph $G$ of treewidth $k$ such that in every tree-decomposition with width at most $ck$, some vertex $v\in V(G)$ has spread strictly more than $c^{\prime}(d(v)+1)$.
Wood also proposed a lower bound example for Problem 1, conjecturing there are no constants $\varepsilon,c>0$ such that $(n\times n)$-grid has a tree-decomposition of width at most $(2-\varepsilon)n$ and spread at most $c$. This conjecture is false and in fact for any $m\geq n\geq 1$ and $\varepsilon>0$, the $(n\times m)$-grid admits a tree-decomposition of width at most $(1+\varepsilon)n$ in which each vertex has spread at most $O(1/\varepsilon)$ (see Lemma 3.1).
When storing a tree-decomposition $(T,(B_{x})_{x\in V(T)})$ of a graph $G$, the memory required depends on $\sum_{x\in V(T)}|B_{x}|$. Our second parameter of interest is the average spread $\frac{1}{|V(G)|}\sum_{x\in V(T)}|B_{x}|$. Regarding this, Wood [24] raised the following problem.
Problem 1.4.
What is the infimum of the $c^{\prime}\in\mathbb{R}$ such that for some $c$ every graph $G$ has a tree-decomposition of width at most $c\operatorname{tw}(G)$ and average spread at most $c^{\prime}$?
Wood already proved the upper bound 3. We show that the answer to this problem is 1.
Theorem 1.5.
Let $c^{\prime}>1$. Then there is a constant $c>0$ such that every graph $G$ has a tree-decomposition of width at most $c\operatorname{tw}(G)$ and average spread at most $c^{\prime}$.
Related work: domino treewidth, tree-partition-width and persistence
In the literature, a special case of tree decompositions with bounded spread was studied, namely those where each vertex has spread at most two. Ding and Oporowski [11] showed that each graph admits such a tree decomposition of width at most $O(\text{tw}(G)^{2}\Delta(G)^{3})$, where $\operatorname{tw}(G)$ and $\Delta(G)$ denote the treewidth and maximum degree of the graph $G$ respectively. Their investigations started from a question of Bienstock [1], who asked whether it is possible for each graph $G$ to find a tree decomposition $(T,(B_{t})_{t\in V(T)})$ in which $T$ is a binary tree and the width as well as the maximum diameter of $T_{v}$ (the subtree of $T$ induced on the bags containing $v$, for $v\in V(G)$), are bounded by a constant depending only on $\operatorname{tw}(G)$ and $\Delta(G)$. Ding and Oporowski [11] answered this question positively in a stronger fashion: their upper bound on the width only depends on $\operatorname{tw}(G)$ and the spread $|V(T_{v})|$ (rather than the diameter) is upper bounded ‘locally’ in terms of the vertex degree $d_{G}(v)$ and $\operatorname{tw}(G)$.
Bodlaender and Engelfriet [6] first used the term domino treewidth for the optimal width of a tree decomposition in which each vertex has spread 2. We write $\operatorname{dtw}(G)$ for the domino treewidth of $G$. Bodlaender [4] improved the upper bound on domino treewidth to $\operatorname{dtw}(G)=O(\operatorname{tw}(G)\Delta(G)^{2})$ and gave examples of graphs $G$ with $\operatorname{dtw}(G)=\Omega(\operatorname{tw}(G)\Delta(G))$.
A tree-partition of a graph $G$ is a partition of $V(G)$ into (disjoint) bags $(B_{i})_{i\in V(T)}$, where $T$ is a tree, such that $uv\in V(G)$ implies that the bags of $u$ and $v$ are the same or adjacent in $T$. The width is the size of the largest bag, and the tree-partition-width $\operatorname{tpw}(G)$ of $G$ is found by taking the minimum width over all tree-partitions of $G$. The notion was first introduced by Seese [22] under the name strong treewidth. {observation} For any graph $G$, $\operatorname{tpw}(G)-1\leq\operatorname{dtw}(G)\leq(\Delta+1)\operatorname{tpw}(G)$. The upper bound follows because after rooting the tree $T$ of a tree-partition of $G$, we may add the contents of each bag to its parent bag to obtain a domino tree decomposition of at most $\Delta+1$ times the width. For the lower bound, note that after rooting a domino tree decomposition, we can remove each vertex from a bag whenever it is not the bag closest to the root containing it. This yields a tree-partition, since for any edge $uv$, at least one of the two is still present in the highest bag they originally shared. (The construction was given by Bodlaender and Engelfriet [6].)
It was first shown that $\operatorname{tpw}(G)=24(\operatorname{tw}(G)+1)\Delta(G)$ by an anonymous referee of [11] and the constant was later improved (to approximately $17.5$) by Wood [23]. The lower bound $2\operatorname{tpw}(G)\geq\operatorname{tw}(G)+1$ was shown by Seese [22]. Tree-partitions for which the underlying tree has bounded maximum degree have been studied as well by Distel and Wood [12].
Bodlaender, Groenland and Jacob [8] showed that Domino Treewidth with the domino treewidth as parameter is complete for the class XALP, i.e., deciding if a given graph has a tree decomposition of width at most $k$ and spread at most $2$ is XALP-complete. They proved a similar result for Tree-Partition-Width and provided an FPT-approximation algorithm for this as well.
Downey and McCartin [13, 14] also studied the notion of spread for path decompositions under the name persistence. They motivated this notion as being a more natural notion for a ‘path-like’ graph, particularly for online computational problems. They showed amongst others that deciding if a given graph has a path decomposition of given width and persistence is W[$t$]-hard for all $t$ via a reduction from Bandwidth, with the sum of the width and persistence as parameter. This reduction can be easily turned into an XNLP-hardness proof, using the recent result by Bodlaender, Groenland, Nederlof and Swennenhuis [9] that Bandwidth is XNLP-complete. An XNLP-membership proof can be done in the same fashion as the XNLP-membership of Bandwidth [9]. Thus Bounded Persistence Pathwidth with sum of width and persistence as parameter is XNLP-complete.
2 Preliminaries
We use the short-cut notation $n]=\{1,\dots,n\}$. A tree decomposition of a graph $G$ is a pair $(T,(B_{t})_{t\in V(T)})$ with $T$ a tree and bags $B_{t}\subseteq V(G)$ for each $t\in V(T)$ such that
  • $\bigcup_{t\in V(T)}B_{t}=V(G)$,
  • for all edges $vw\in E(G)$, there is an $t$ with $v,w\in B_{t}$, and
  • for all $v$, the nodes $\{t\in V(T)\penalty 10000\ |\penalty 10000\ v\in B_{t}\}$ form a subtree (denoted $T_{v}$) of $T$.
The width of the decomposition is $\max_{t\in V(T)}|B_{t}|-1$, and the treewidth of a graph $G$ is the minimum width over all tree decompositions of $G$. A path decomposition is a tree decomposition $(T,(B_{t})_{t\in V(T)})$ for which $T$ is a path and the pathwidth is the minimum width over all path decompositions of $G$. The spread of a vertex $v$ in a tree-decomposition $(T,(B_{t})_{t\in V(T)})$ is the number of bags $v$ is in: $|\{t\in V(T):v\in B_{t}\}|$.
The following lemma is standard (see e.g. [24, Corollary 9]).
Lemma 2.1 (Folklore).
Let $(T,(B_{t})_{t\in V(T)})$ be a tree decomposition of $G$. Let $w:V(G)\to\mathbb{R}_{\geq 0}$ be a weight function. Then there is $t\in V(T)$ such that for each component $C$ of $G-B_{t}$, we find that $w(C)\leq w(V(G))/2$.
3 Grid graphs and extensions
The $(n\times m)$-grid is the graph on vertex set $n]\times[m]$ and edge set
$\{(i,j)(i+1,j):i\in[n-1]\}\cup\{(i,j)(i,j+1):j\in[m-1]\}.$
We first provide a near-optimal width tree decomposition for grids with constant spread.
Lemma 3.1.
Let $c>1$ and let $n,m\in\mathbb{N}$. Then the $(n\times m)$-grid has a path decomposition of width $cn$ and spread at most $\lceil 1/(c-1)\rceil+1$.
We provide a ‘proof by picture’ below and included a formal proof in the appendix.
The picture above shows four consecutive bags for $c=1+1/3$ via the coloured regions. Each column denotes a single column of the $(n\times m)$-grid, whereas any row in the picture corresponds to the union of at most $\lceil n/3\rceil$ consecutive rows of the grid.
Let $D_{n,m}$ denote the graph obtained from the $(n\times m)$-grid by adding the same diagonal edge in each square, that is, we add the edges
$\{(i,j)(i+1,j+1):i\in[n-1],j\in[m-1]\}.$
When $n=m$, this can also be seen as the dual graph of the hexagonal grid.
We also define the east, west, north and south sets as
$E$ $=\{(i,1):i\in[n]\},$
$W$ $=\{(i,m):i\in[n]\},$
$N$ $=\{(1,j):j\in[m]\},$
$S$ $=\{(n,j):j\in[m]\}.$
A picture is given below.
NNSSEEWW
The following result is well-known (see e.g. [15] for the case $n=m$ and the introduction of [17] for a sketch of the result below).
Theorem 3.2 (Variant of Hex theorem).
Let $n,m\geq 1$ be integers. For any two sets $H$ and $V$ covering $V(D_{n,m})$, either there is a path in $D_{n,m}[H]$ from $E$ to $W$ or there is a path in $D_{n,m}[V]$ from $N$ to $S$.
We call such a path from $N$ to $S$ in the theorem statement above a north-south-path. From this result, we obtain the following corollary.
Corollary 3.3.
If $S\subseteq V(D_{n,m})$ does not contain a north-south-path, then $D_{n,m}-S$ has a connected component of size at least n(m|S|)n(m-|S|).
Proof 3.4.
We apply Theorem 3.2 with $V=S$ and $H=V(D_{n,m})\setminus S$ to find that there is a connected component $C$ of $D_{n,m}-S$ which contains a vertex on every column of $D_{n,m}$. Since the set $S$ contains vertices in at most $|S|$ columns, for any other column $j$, the vertices $\{(i,j):i\in[n]\}$ in this column will all be in $C$. Hence $|C|\geq n(m-|S|)$.
Here it becomes clear why it is convenient to move to unbalanced grids: when $m$ is much larger than $n$, any ‘balanced’ separator now must contain a north-south-path.
Let $m=n^{8}$ and split $[m]$ into $b=n^{4}$ consecutive intervals $J_{1},\dots,J_{b}$ of length $m^{\prime}=n^{4}$. For each $i\in[b]$, we choose $y_{i}\in J_{i}$. Let $D_{n,m}^{+}$ be the graph obtained from $D_{n,m}$ by adding a vertex $v_{0}$ with neighbourhood
$\{(1,y_{1}),(1,y_{n^{2}+1}),\dots,(1,y_{b-n^{2}+1})\}.$
In particular, $d(v_{0})=b/n^{2}=n^{2}$. We call $v_{0}$ the special vertex. A visualisation is given below.
[…]
Lemma 3.5.
For $n$ sufficiently large, with $m=n^{8}$, for any tree decomposition $(T,(B_{x})_{x\in V(T)})$ of $D_{n,m}^{+}$ of width 2n22n-2, the special vertex $v_{0}$ has spread at least $\frac{1}{2}n^{2}(d(v_{0})+1)$ in the tree decomposition.
Proof 3.6.
Let $(T,(B_{x})_{x\in V(T)})$ be a tree decomposition of $D_{n,m}^{+}$ of width 2n22n-2. Let $v_{0}$ denote the special vertex of $D_{n,m}^{+}$.
For $a\in[b]$, let $w_{a}:V(D_{n,m}^{+})\to\{0,1\}$ be defined by $w_{a}((i,j))=1$ if $j\in J_{a}$ and 0 otherwise. For $a\in[b]$, applying Lemma 2.1 with $w_{a}$ as weighting, provides $x(a)\in V(T)$ such that each component $C$ of $D_{n,m}^{+}-B_{x(a)}$ satisfies $w_{a}(C)\leq w_{a}(V(G))/2$. Since
$n(m^{\prime}-|B_{x(a)}|)\geq n(n^{4}-(2-\varepsilon)n)>n^{5}/2=nm^{\prime}/2,$
Corollary 3.3 applied to the copy of $D_{n,m^{\prime}}$ induced on $n]\times J_{a}$, shows that $B_{x(a)}$ must contain a north-south-path which stays within $n]\times J_{a}$. Let $X_{a}\subseteq[n]\times J_{a}$ denote $n$ vertices in $B_{x(a)}$ chosen so that each $i\in[n]$ occurs exactly once as first coordinate.
Figure 1: The figure shows a potential north-south path (in red) in $J_{a}$ and highlights the vertices chosen in $X_{a}$ (in orange).
A visualisation is given in Figure 1.
We next argue that there is a subpath $P$ of $T$ which contains the nodes $x(1),\dots,x(b)$ in this particular order. There are two alternatives to exclude.
  • Suppose that for some $1\leq i_{1}<i_{2}<i_{3}\leq b$, the node $x(i_{3})$ lies on the path between $x(i_{1})$ and $x(i_{2})$ in $T$. Within $D_{n,m}$, there are $n$ vertex disjoint paths from $X_{i_{1}}$ to $X_{i_{2}}$ that do not pass through $X_{i_{3}}$. In particular, $B_{x(i_{3})}$ now must contain one vertex for each such path, plus the $n$ vertices in $X_{i_{3}}$. But then $|B_{x(i_{3})}|\geq 2n$, contradicting the fact that all bags have size at most 2n12n-1.
    A similar argument applies to exclude the possibility that $x(i_{1})$ lies in between $x(i_{2})$ and $x(i_{3})$.
  • Suppose that for some $1\leq i_{1}<i_{2}<i_{3}\leq b$, there is a node $y\in V(T)$ which has three internally vertex-disjoint paths to $x(i_{1}),x(i_{2})$ and $x(i_{3})$. Then $|B_{y}|\geq 2n$ (contradicting the width again) since it must contain at least one vertex of each of the following $2n$ vertex-disjoint paths: the $n$ paths from $X_{i_{1}}$ to $X_{i_{2}}$ which keep their first coordinate fixed and the $n$ paths from
    $X_{i_{2}}+(0,1)=\{(x,y+1):(x,y)\in X_{i_{2}}\}$
    to $X_{i_{3}}$ which keep their first coordinate fixed.
So $T$ has a subpath $P$ containing the nodes $x(1),\dots,x(b)$ in this particular order. Next, we show that
$v_{0}\in B_{x(n\cdot n^{2}+1)}\cap B_{x((n+1)n^{2}+1)}\cap\dots\cap B_{x(b-n\cdot n^{2}+1)}.$ (1)
Suppose first towards a contradiction that $v_{0}\not\in B_{x(a)}$ for all $a<j$ yet $v_{0}\in B_{x(j)}$ for some $j\in\{n\cdot n^{2}+1,\dots,b\}$. For $\ell\in\{0,1,\dots,n-1\}$, we choose a vertex $x_{\ell}\in X_{\ell n^{2}+1}$ and define a path $P_{\ell}$ within $J_{\ell n^{2}+1}\cup\{v_{0}\}$ running from $v_{0}$ to $x_{\ell}$ via the neighbour $(1,y_{\ell n^{2}+1})$ of $v_{0}$. The paths $P_{\ell}$ intersect only in $v_{0}$ and are disjoint from $X_{j}$. Recall that the vertices in $X_{a}$ are by definition in the bag $B_{x(a)}$. Since $v_{0}\not\in B_{x(a)}$ for all $a<j$ and $v_{0}\in B_{x(j)}$, for each $\ell\in\{0,1,\dots,n-1\}$, $B_{x(j)}$ must either contain $x_{\ell}$, or $B_{x(j-1)}\cap B_{x(j)}$ is a separator between $x_{\ell}$ and $v_{0}$. Either way, $B_{x(j)}$ must contain a vertex of $P_{\ell}\setminus\{v_{0}\}$. Since the $n$ vertices from $X_{j}$ are also present in $B_{x(j)}$, using the disjointness constraints, we obtain $2n$ different vertices in $B_{x(j)}$, contradicting the upper bound on the width. A similar argument can be used with the neighbours $(1,y_{b-n^{2}+1}),\dots,(1,y_{b-n\cdot n^{2}+1})$ to exclude the possibility that $v_{0}\not\in B_{x(a)}$ for all $a>j$ yet $v_{0}\in B_{x(j)}$, for all $j\in\{1,\dots,b-n\cdot n^{2}+1\}$.
We showed above that if $v_{0}\in B_{x(j)}$ for some $j\in[b]$, then it is also in $B_{x(a)}$ for all $a\in\{j,\dots,b-n\cdot n^{2}+1\}\cup\{n\cdot n^{2}+1,\dots,j\}$. The remaining case in which (1) fails is when $v_{0}\not\in B_{x(j)}$ for all $j\in[b]$. Then there is a unique $j$ such that $x(j)$ is closest among $x(1),\dots,x(b)$ to a bag containing $v_{0}$. The same argument as above can now be applied, since $B_{x(j-1)}\cap B_{x(j)}$ (or $B_{x(j)\cap B_{x(j+1)}}$) will still form a separator in this case between $v_{0}$ and $x_{\ell}$.
In conclusion, we established that $v_{0}$ is present in at least $b-2(n+1)n^{2}=n^{4}-2(n+1)n^{2}$ bags while its degree is $b/n^{2}=n^{2}$. For $n$ sufficiently large, the spread of $v_{0}$ is at least $n^{4}-2(n+1)n^{2}>\frac{1}{2}n^{2}(n^{2}+1)=\frac{1}{2}n^{2}d(v)$.
The lemma above implies Theorem 1.3 for all $c^{\prime}<2$ since $D_{n,m}^{+}$ has treewidth at most $n+2$ (since adding $v_{0}$ to all bags in a tree decomposition of $D_{n,m}$ provides a tree decomposition for $D_{n,m}^{+}$).
4 Well-behaved spread using thrice the width
In this section we prove Theorem 1.2 which is immediately implied by the result below.
Theorem 4.1.
Let $b\geq 1$ be an integer. Every graph $G$ admits a tree decomposition of width at most $(3+8/b)(\textup{tw}(G)+1)$ for which each vertex $v\in V(G)$ is in at most $2b(d(v)+1)$ bags.
A key idea in the best bound of 14 from Wood [24] is the definition of a slick tree decomposition $(T,(B_{x})_{x\in V(T)})$, in which the tree $T$ is rooted and the following property holds: if a vertex $v\in V(G)$ is in $B_{x}\cap B_{y}$ where $x$ is the parent of $y$, then $v$ has a neighbour in $B_{y}\setminus B_{x}$. This property automatically ensures that each vertex has spread at most $d(v)+1$.
Using separators, a rooted tree decomposition can be constructed using induction on the number of vertices in the graph. After the first separator $S$ is obtained, a recursive call can be placed on the components of GSG-S; the aim is then to glue together the tree decompositions obtained via recursion using a bag containing $S$. At the next recursion depth, we are given a special set $S$ that needs to become part of the root bag (thinking ahead of our gluing operation), and we find a ‘balanced’ separator $X$ for $S$, passing on $X$ and parts of $S$ to the next recursion depth as special set. Such an approximation scheme for treewidth has been used since the 1990s [7, 20], see [5] for a recent overview.
Wood [24] cleverly noticed that in this routine, before passing a special set $S^{\prime}$ to a recursive call, we can also replace it by a set $S$ where we add some neighbours of vertices in $S$ in order to ensure the tree decomposition will become slick. In order to keep the bag size smaller compared to their approach, we will only add neighbours for a ‘negligible’ fraction of the vertices, while ensuring all vertices have a neighbour added once-in-a-while to upper bound the spread.
To facilitate the reasoning about the spread of the vertices, we introduce the notion of a $b$-marking for a rooted tree decomposition $(T,(B_{t}))$ of $G$. This is a partial function $m:V(G)\times V(T)\to\{1,\dots,b\}$ such that:
  • $m(v,t)$ is defined if and only if $v\in B_{t}$;
  • if $v$ is in $B_{c}\cap B_{p}$ for $c$ the child of $p$ in $T$, then either $m(v,c)<m(v,p)$ or $v$ has a neighbour in $B_{c}$ which it does not have in $B_{p}$;
  • if $v\in B_{\ell}$ and $\ell$ is a leaf of $T$ (not the root) then $v$ has a neighbour in $B_{\ell}$ which it does not have in the parent bag of $\ell$.
Note that this last property can always be ensured, since we only need to put vertices in leaf bags if they have a new neighbour in this bag (else we can just remove them from the bag).
The use of this notion is via the following observation.
Lemma 4.2.
If a rooted tree decomposition admits a $b$-marking, then each vertex $v$ is in at most $2b(d(v)+1)$ bags.
Proof 4.3.
Let $(T,(B_{t})_{t\in V(T)})$ be a tree decomposition of $G$ with $b$-marking $m$. Let $v\in V(G)$. Let $T_{v}$ be the subtree of $T$ corresponding to the bags containing $v$. We call a node of $T^{\prime}$ happy if there is a neighbour of $v$ in $B_{t}$ which was not in $B_{p}$ for $p$ the parent of $t$. (The root is also happy.) Note that:
  • each leaf node of $T_{v}$ is happy;
  • any path of length at least $b$ with only degree 2 vertices has at least one happy vertex.
The second property holds since such a path $t_{1},\dots,t_{b}$ must pass from parent to child in each step and therefore the marking goes down by 1, or a happy node occurs, by definition. We use this to show that the fraction of nodes in $V(T_{v})$ that is happy is at least $1/(2b)$. Since the number of non-root happy nodes is at most the number of neighbours of $v$, this shows that $|V(T_{v})|\leq 2b(d(v)+1)$.
The remainder of the proof is straightforward. We replace $T_{v}$ by a tree $T^{\prime}$ by contracting greedily any path of length at least $b$ with only degree 2 vertices. Note that among the nodes we removed this way, a fraction of $1/b$ is happy, so it suffices to show a fraction of $1/(2b)$ in $T^{\prime}$ is happy. In $T^{\prime}$, at most a ((b1)/b)((b-1)/b)-fraction of the vertices have degree 2; half of the remaining vertices must have degree 1 (by the handshake lemma) and those are all happy. This shows that at least a $1/(2b)$-fraction of the vertices in $T^{\prime}$ and so in $T_{v}$ is happy, as desired.
The constant 22 in the lemma above is probably not tight, but this is irrelevant for our aims.
To prove Theorem 4.1, we prove a slightly stronger claim by induction on $n$ below.
Lemma 4.4.
Let $b\geq 4$ and $k\geq 2$ be integers. Let $G$ be a graph on $n$ vertices with treewidth k1k-1. Let $S\subseteq V(G)$ be given with
$k+1\leq|S|\leq(2+8/b)k$
together with $m^{\prime}:S\to\{1,\dots,b\}$ such that at most $(1/b)(2+8/b)k$ vertices in $S$ are marked $i$ by $m^{\prime}$ for all $i\in[b-1]$. Then $G$ admits a tree decomposition $(T,(B_{x})_{x\in V(T)})$ such that
  • $|B_{x}|\leq(3+8/b)k$ for all $x\in V(T)$,
  • $T$ is rooted in $r$ and $S\subseteq B_{r}$,
  • there is a $b$-marking $m:V(G)\times V(T)\to[b]$ for which $m(s,r)=m^{\prime}(s)$ for all $s\in S$.
Proof 4.5.
If $|V(G)|\leq(3+8/b)k$ then the lemma holds since we may put all vertices in the same bag.
Suppose that $|V(G)|>(3+8/b)k$ and that the lemma has been shown already for graphs on fewer vertices. Let $S$ be given as in the lemma and give all vertices in $S$ weight $1$ and the remaining vertices weight 0. We apply Lemma 2.1 with this weight function to find a $t$ for which for each component $C$ of $G-B_{t}$, we find $|C\cap S|\leq|S|/2$.
Let $X=B_{t}$. For each component $C$ of GXG-X, we add vertices to a set $S_{C}$ and define a marking $m_{C}^{\prime}:S_{C}\to[b]$ as follows:
  • For all vertices $v$ in $S$ marked 1 by $m^{\prime}$ (which may be in $X$) that have a neighbour in $C\setminus S$, we add $v$ and a neighbour of $v$ in $C\setminus S$ to $S_{C}$, give $v$ mark $b$ and also give the neighbour mark $b$.
  • All other vertices $s$ in $S\cap C$ that have a neighbour in $C\setminus S$, we add with mark $m^{\prime}(s)-1$. (Here $m^{\prime}$ was given in the lemma statement.)
  • All vertices $x$ in $X\setminus S$ with a neighbour in $C\setminus S$, we add with mark b1b-1.
In particular, vertices in $X$ or $S$ without neighbours in $C\setminus S$ are not added to $S_{C}$. Note that
$|S_{C}|\leq k+|S|/2+(1/b)(2+8/b)k\leq(1+1+4/b+2/b+8/b^{2})\leq(2+8/b)k$
for $b\geq 4$. Here we used that at most $(1/b)(2+8/b)k$ vertices had mark 1 via $m^{\prime}$.
We wish to apply the induction hypothesis on $G_{C}=G[(C\setminus S)\cup S_{C}]$ with special set $S_{C}$ and the marking $m^{\prime}_{C}$ constructed above. If $|V(G_{C})|\leq 3k$, then this is not needed and we may simply define $R_{C}=S_{C}$. Otherwise, we can ensure that $|S_{C}|\geq k+1$; if needed, we can add vertices from $C$ to $S_{C}$ and mark these $b$ by $m_{C}^{\prime}$.
In order to ensure the marking $m^{\prime}_{C}$ for $S_{C}$ satisfies the assumptions in the induction hypothesis/lemma statement, we can decrease some of the marks if needed. By definition, at this point at most $(1/b)(2+8/b)k$ vertices in $S_{C}$ are marked $i$ for some $i\in[b-2]$, since only vertices that were previously marked $i+1$ in $S$ by $m^{\prime}$ can receive mark $i$ for $i\in[b-1]$ by $m_{C}^{\prime}$. Moreover, at most
$|S\cap C|+|X|\leq|S|/2+k\leq(1-1/b)(2+8/b)k$
vertices can ever receive a mark $\leq b-1$. This means that if too many vertices received mark b1b-1 at this point, we can safely decrease some of their marks (‘increasing their priority’).
We are now ready to apply the induction hypothesis on $G_{C}=G[(C\setminus S)\cup S_{C}]$ with special set $S_{C}$ and the marking $m^{\prime}_{C}$ adjusted as above (if needed). This yields a tree decomposition $T_{C}$ with root bag $R_{C}$ containing $S_{C}$ for each component $C$ together with a $b$-marking $m_{C}$.
We make one additional bag $B_{r}=S\cup X$ and make it adjacent to the various $R_{C}$ bags. This creates a tree decomposition which we root in $B_{r}$. We let the marking $m$ be defined as by
  • $m(s,r)=m^{\prime}(s)$ for all $s\in S$;
  • $m(x,r)=b$ for all $x\in X\setminus S$;
  • $m(v,t)=m_{C}(v,t)$ for $t\in V(T_{C})$ and $v\in B_{t}$.
In order to show this is a $b$-marking, we only need to consider what happens to $v\in B_{p}\cap B_{c}$ where $B_{p}=B_{r}=S\cup X$ is the root bag and $B_{c}=R_{C}$ for some component $C$. If $v\in X\setminus S$, then
$m(v,c)=m_{C}(v,c)=m^{\prime}_{C}(v)\leq b-1=m(v,p)-1$
as desired. If $v\in S$ then
$m(v,c)=m_{C}(v,c)=m^{\prime}_{C}(v)=m^{\prime}(v)-1=m(v,p),$
unless $m^{\prime}(v)=1$ in which case a neighbour of $v$ in $C$ which was not in $B_{r}$ will have been added. This shows that $m$ is a $b$-marking, as desired.
Proof 4.6 (Proof of Theorem 4.1).
Let $b\geq 1$ be an integer and let $G$ be a graph and let $k=\textup{tw}(G)+1$. The statement is immediately true when $|V(G)|\leq k$. Otherwise, we let $S$ contain $k+1$ vertices of $G$ and set $m^{\prime}(s)=b$ for all $s\in S$. Lemma 4.4 now provides a tree decomposition of width at most $(3+8/b)k=(3+8/b)(\text{tw}(G)+1)$ that admits a $b$-marking. By Lemma 4.2 each vertex $v$ has spread at most $2b(d(v)+1)$ in this.
5 Proof of Theorem 1.5
We will use the following auxiliary result.
Lemma 5.1 (Lemma 22 in [24]).
For any integer $k\geq 2$, every rooted tree $T$ with $|V(T)|\geq k$ with root $r$ has a sequence $(T_{1},\dots,T_{m})$ of pairwise edge-disjoint rooted subtrees of $T$ such that:
  • $T=T_{1}\cup\dots\cup T_{m}$;
  • $r\in V(T_{1})$ and for $i\in\{2,\dots,m\}$, if $r_{i}$ is the root of $T_{i}$ then $V(T_{i})\cap V(T_{1}\cup\dots\cup T_{i-1})=\{r_{i}\}$;
  • $|V(T_{i})|\in\{k,\dots,2k-2\}$ for each $i\in\{1,\dots,m\}$.
In fact, our proof of Theorem 1.5 will follow an approach from Wood [24], for which we apply the lemma with a different parameter. Only the analysis of the spread in this construction is new.
Proof 5.2 (Proof of Theorem 1.5).
Let $c^{\prime}>1$. Choose an integer $t\geq 1$ such that
$t/(t-1)\leq c^{\prime}.$
Let $c=1+2t$. We will prove that every graph $G$ has a tree-decomposition $(T,(B_{t})_{t\in V(G)})$ of width at most $c\operatorname{tw}(G)$ and average spread
$\frac{1}{|V(G)|}\sum_{x\in V(G)}|B_{x}|\leq c^{\prime}.$
Let $G$ be a graph of treewidth $w$. Then there is a tree decomposition $(T,(B_{x})_{x\in V(T)})$ of $G$ with width $w$, which we can root such that if $x$ is the parent of $y$ in $T$, then $|B_{y}\setminus B_{x}|=1$, that is, each bag contributes exactly one new vertex. We apply Lemma 5.1 to $T$ with $k=t\cdot w$ to obtain $(T_{1},\dots,T_{m})$. Let
$B_{1}^{\prime}=\cup_{s\in V(T_{1})}B_{s}$
and for $i\in\{2,\dots,m\}$, let
$B_{i}^{\prime}=\cup_{s\in V(T_{i})\setminus\{r_{i}\}}B_{s}.$
In [24, Lemma 23] it is shown that such a quotient yields a tree decomposition $(F,(B_{x}^{\prime})_{x\in V(F)})$ of $G$, for any tree $F$ with vertex-set $\{1,\dots,m\}$, rooted at vertex $1$, where for $i\in\{2,...,m\}$, the parent of $i$ is any number $j\in\{1,\dots,i-1\}$ such that $r_{i}\in V(T_{j})$.
Note that for any $i\in[k]$,
$|B_{i}^{\prime}|\leq w+|V(T_{i})|\leq w+2k=(1+2t)w=c\cdot\text{tw}(G),$
since adjacent bags introduce at most one new vertex. Moreover,
$|B_{i}^{\prime}|\geq|V(T_{i})|\geq k=t\cdot w.$
since each bag introduces at least one new vertex.
We say that vertex $v$ is new in bag $B_{i}^{\prime}$ if $i$ is the smallest $j\in[m]$ for which $v\in B_{j}^{\prime}$.
Claim 1.
For all $i\in[m]$, $B_{i}^{\prime}$ has at least $(1-1/t)|B_{i}^{\prime}|$ new vertices.
Proof 5.3.
The new vertices in $B_{i}^{\prime}$ are exactly those in $B_{i}^{\prime}\setminus B_{r_{i}}$. Indeed, $T_{i}$ only intersects $T_{1},\dots,T_{i-1}$ in $r_{i}$ and since $(T,(B_{s})_{s\in V(T)})$ is a tree decomposition,
$\left(\cup_{s\in V(T_{i})\setminus\{r_{i}\}}B_{s}\right)\cap\left(\cup_{s\in V(T_{1})\cup\dots\cup V(T_{i-1})}B_{s}\right)\subseteq B_{r_{i}}.$
Using that for each $s\in V(T_{i})\setminus\{r_{i}\}$, there is at least one vertex in $B_{s}$ that was not in the bag of its parent, we find that
$|B_{r_{i}}\cap B_{i}^{\prime}|/|B_{i}^{\prime}|\leq w/(w+|V(T_{i})|-1)\leq w/(w+wt-1)\leq 1/t,$
and so there are at least $(1-1/t)|B_{i}^{\prime}|$ new vertices in $B_{i}^{\prime}$.
We now show how the claim above implies that the average spread $\sum_{x\in V(F)}|B_{x}^{\prime}|/|V(G)|$ of the tree decomposition is at most
t/(t1)=1+1/(t1)t/(t-1)=1+1/(t-1)
For $v\in V(G)$, let $i_{v}$ denote the (unique) index $i$ such that $v\in B_{i}^{\prime}$ is new. Then
$|V(G)|=|\{(v,B_{i_{v}}^{\prime}):v\in V(G)\}|=\sum_{i=1}^{m}|\{(v,B_{i}^{\prime}):i=i_{v},v\in V(G)\}|.$
By the claim, for each $i$, since the summand is exactly the number of new vertices,
$|\{(v,B_{i}^{\prime}):i=i_{v},v\in V(G)\}|\geq(1-1/t)|B_{i}^{\prime}|.$
This proves
$|V(G)|\geq\sum_{i=1}^{m}|B_{i}^{\prime}|(1-1/t).$
Rearranging shows that the average spread
$\sum_{i=1}^{m}|B_{i}^{\prime}|/|V(G)|\leq 1/(1-1/t)=1/((t-1)/t)=t/(t-1).$
Hence we provided the desired tree decomposition.
6 Conclusion
In this paper we increased our understanding of the trade-off between width and spread for tree decompositions. We discuss the algorithmic aspects and two conjectures below.
Algorithmic aspects.
The proof of Theorem 4.1 can easily be made algorithmic. Given a $c$-approximation algorithm for finding balanced separators, for any integer $b\geq 1$ a tree decomposition can be computed for any graph $G$ of width at most $c(3+8/b)(\operatorname{tw}(G)+1)$ for which each vertex $v\in V(G)$ is in at most $2b(d(v)+1)$ bags, at only polynomial overhead cost. Similarly, the proof of Theorem 1.5 can also be made algorithmic, showing that for any $c^{\prime}>1$, there is a constant $c>0$ and a polynomial time algorithm which computes for any graph $G$ a tree-decomposition of width at most $c\operatorname{tw}(G)$ and average spread at most $c^{\prime}$.
Although notions related to spread have been motivated several times for potential algorithmic use, we are not currently aware of any algorithms that explicitly use tree decompositions with small spread and width. A recent application of a result on tree decompositions with additional properties, namely when the tree of the decomposition is relatively small, is in the construction of universal graphs [19].
Conjecture for Problem 1
In terms of the solution to Problem 1, we expect that the answer is 3. The following example could perhaps be used to provide a lower bound which matches our upper bound from Theorem 1.2. In Section 3 we defined $D_{n,m}^{+}$ for $m=n^{8}$, which is obtained from the $(n\times m)$-grid by consistently adding one diagonal per square and adding a special vertex with neighbours spread among the first row. Let $H$ be the graph obtained from $n^{2}$ copies of $D_{n,m}^{+}$, which are glued in a sequence by gluing the ‘west boundary’ of the $i$th copy on the ‘east boundary’ of the $(i+1)$th, with a vertex $v_{0}$ added which is adjacent to $n^{2}-1$ vertices that lie on the first row of these ‘glued boundaries’. Let $v_{0}^{(i)}$ denote the special vertex of the $i$th copy of $D_{n,m}^{+}$. Suppose that a tree decomposition of $H$ has width <3n<3n. We expect that if any bag contains two north-south paths that have many neighbours of $v_{0}^{(i)}$ in between them, then the spread of $v_{0}^{(i)}$ is forced to be large. However, when not allowing such a bag, it is difficult to ensure that the bags containing the neighbours of $v_{0}$ are all ‘close to each other’ and thereby the spread of $v_{0}$ becomes large. We did not see a nice way of formalising this intuition and so we leave open what the exact answer is.
Spread versus width in the grid.
Wood [24] also conjectures that every tree-decomposition of the $(n\times n)$-grid with width $n$, has some vertex with spread $\Omega(n)$. We propose the following stronger conjecture, which would be best possible up to the multiplicative constant.
Conjecture 6.1.
There is a constant $\delta>0$ such that for all integers $m\geq n\geq 1$, every tree-decomposition of the $(n\times m)$-grid with width $n+a$, has a vertex with spread at least $\delta(n/a)$.
We sketch why this conjecture holds in the case that $m$ is much larger than $n$. We may assume that $a\leq\varepsilon n$ (since we may choose $\delta$ and each vertex has spread at least 1). Using Lemma 2.1, we can find a bag $B_{t}$ that is a balanced separator. Using a similar argument to the proof of Corollary 3.3, we find that $\{(i,x_{i}):i\in[n]\}\subseteq B_{t}$ for some choice of $x_{1},\dots,x_{n}\in[m/3,2m/3]$. Let $I\subseteq[n]$ be the set of $i\in[n]$ for which $(i,x_{i})$ is the unique vertex of the form $(i,j)$ in the bag $B_{t}$. Then $|I|\geq n-a$. For $m$ much larger than $n$, all vertices in $\{(i,x_{i}+1):i\in I\}$ then must also be in the same connected component, and so there is a neighbour $t_{2}$ of $t$ with $\{(i,x_{i}):i\in I\}\subseteq B_{t_{2}}$. We can then find a neighbour $t_{3}$ of $t_{2}$ that loses at most $a$ more vertices from the set $I$, and repeat this $n/a$ before $I$ is empty. This shows that some vertex of the form $(i,x_{i})$ is in at least $n/a$ bags.
For the $(n\times n)$-grid, we find it likely that a similar situation occurs, but in this case the separator structure becomes more complicated. Although it is no longer the case that each balanced separator contains a north-south path, we remark that a weaker analogue of this still holds.
Lemma 6.2 (Bodlaender [3, Claim 88a]).
Let $(T,(B_{x})_{x\in V(T)})$ be a tree decomposition of an $(n\times n)$-grid of treewidth $n$. Then there is a $t\in V(T)$, such that at least one of the following two holds:
  1. 1.
    $B_{t}$ contains a vertex from each row of the grid.
  2. 2.
    $B_{t}$ contains a vertex from each column of the grid.
We remark that for the $(n\times n)$-grid, it is already highly non-trivial to prove the rankwidth is n1n-1 (as shown by Jelínek [18]) or that the cliquewidth is $n+1$ (as shown by Golumbic and Rotics [16]).
Appendix A Omitted proof
Recall that Lemma 3.1 states that for any $c>1$, any $(n\times m)$-grid has a path decomposition of width $cn$ and spread at most $\lceil 1/(c-1)\rceil+2$.
Proof A.1 (Proof of Lemma 3.1).
We divide $[n]$ into at most $b=\lceil 1/(c-1)\rceil$ consecutive intervals of size at most $\lceil(c-1)n\rceil$, say $n]=A_{1}\sqcup A_{2}\sqcup\dots\sqcup A_{b}$. Let $A_{[i,j]}=A_{i}\cup A_{i+1}\cup\dots\cup A_{j}$. We claim the following sequence of bags forms the desired path decomposition. The sequence starts with
$(A_{[1,b]}\times\{1\})$ $\cup(A_{1}\times\{2\})$
$(A_{[2,b]}\times\{1\})$ $\cup(A_{[1,2]}\times\{2\})$
$(A_{[3,b]}\times\{1\})$ $\cup(A_{[1,3]}\times\{2\})$
$\vdots$
$(A_{b}\times\{1\})$ $\cup(A_{[1,b]}\times\{2\})$
$(A_{[1,b]}\times\{2\})$ $\cup(A_{1}\times\{3\})$
and continues in this fashion. To be precise, for $i\in[b]$ and $j\in[m-1]$, let
$B_{i,j}=(A_{[i,b]}\times\{j\})\cup(A_{[1,i]}\times\{j+1\})$
where $1,1]=\{1\}$ and $b,b]=\{b\}$. The sequence is
$B_{1,1},B_{2,1},\dots,B_{b,1},B_{1,2},\dots,B_{b,2},\dots,B_{1,m-1},\dots,B_{b,m-1}.$
Note that each bag $B_{i,j}$ has size at most $n+\lceil(c-1)n\rceil=\lceil cn\rceil$, so the width is at most $cn$ as claimed. Each vertex of the $n\times m$ grid is in $A_{i}\times\{j\}$ for exactly one value of $(i,j)$. Each $A_{i}\times\{j\}$ appears in the sequence above at most $b+1$ times. So the spread of each vertex is at most $b+1=\lceil 1/(c-1)\rceil+1$.
Finally, we verify that the sequence of bags forms a path decomposition directly from the definition: the bags containing a vertex form a subpath and every vertex is covered. All edges are of one of the following forms:
$\{(x,j),(x+1,j)\},\penalty 10000\ \{(x,j),(x,j+1)\}.$
Let $i$ be given such that $x\in A_{i}$. Then $(x,j),(x+1,j),(x,j+1)\in B_{i,j}$, so there is an edge containing both types of edges above.